DM Monthly Status Report for February 2017

The DM monthly status report covering February has been posted in Docushare, Collection-5185. For your convenience I am pasting high-level and summary and project status below. Direct link to full report (pdf):

##High-level Summary
The DM Team prepared for the JTM in March, continued to prepare for the arrival of the DM Project Manager in Tucson, continued supporting LSST System Engineering and Operations Planning activities, added features deblending and spectral data processing and improved performance of the LSST Software Stack and completed the technical portion of R13.0, added all-sky catalog search, image cutout display features to the Prototype DAC (PDAC), completed the installation of the Mountain - Base fiber optic cable, prepared for large-scale pre-cursor data production runs at the Archive Center, and developed a draft Service Level Agreement and prototype SLA database.
##Project Status
Selected technical accomplishments:

  • The Architecture Team:
    • Prepared for the Joint Technical Meeting in March, worked on the DM Design, DM System Requirements, and EPO interface documents, and convened a new SuperTask/Butler Working Group.
  • The AP Team:
    • Completed detailed comparison of image differencing algorithms on simulated data.
    • Implemented ability to use different photometric and astrometric catalogs in jointcal.
    • Demonstrated scaling of producers and consumers for the prototype alert distribution system.
  • The DRP Team:
    • Demonstrated outstanding results from the prototype Non-negative Matrix Factorization (NMF) deblender.
    • Produced an early design sketch of “MultiFit”-style multi-epoch fitting system.
    • Tested a prototype spectral data reduction pipeline for the LSST Auxiliary Telescope.
  • The SUIT Team:
    • Continued to add new functions to PDAC deployed at NCSA, allsky catalog search, image cutout display.
    • Bug fixes and improvement in Python API.
  • The DAX Team:
    • Continued support of integration activities at the PDAC via problem investigations and bug fixes.
    • Completed ingest of 20% DR1 synthetic test dataset at IN2P3, to be used in gathering Qserv KPMs.
  • The Net&BS Team:
    • All the fusion splices were completed on the Summits - AURA Gatehouse Network in February and it only remains to terminate the cable at the LSST telescope when we have occupancy of the computer room, which is expected to occur NLT September, 2017.
    • The Telefonica fiber installation from Highway D41 to Gatehouse is complete. There was a final problem with the fiber installation from D41 to the Gatehouse, which was solved by installing three more posts at the crossing point of D41 and the Observatory access road.
    • The first set of trenches and ducts that will be required to bring the new fibers into the compound in La Serena have been completed. There are now ducts available to bring fibers into the compound to the existing data center.
  • The NCSA Team:
    • Held meetings with CC-IN2P3 to discuss plans for coordinated operations, including data movement between sites and a role in disaster recovery.
    • Completed the first round of operational planning discussions with FNAL.
    • Prepared for running HSC datasets through the existing batch production system pipeline.
    • Proposed a design for a header metadata client and discussed it with OCS/EFD staff.
    • Made significant progress detailing operational requirements for database use across the data backbone, including provenance and metadata catalogs.
    • Created a first draft Service Level Agreement in the prototype service management framework.
    • Tuned performance of the verification and compute clusters in response to user feedback.
    • Created a detailed design of the FY17 L1 test system.
  • The SQuaRE Team:
    • Added HSC data to our verification harness, SQuaSH.
    • Completed the technical part of the 13.0 release.
    • Deployed the version dashboard for our LTD-hosted documents.
    • Deployed a simple metrics monitoring microservice that delivers notifications of metric deviations via our chatbot.
    • Deployed a new version of our git-lfs service.
    • Extended our OSX build capacity.